# Recurrence-based Vanishing Point Detection

### v1

@@ 1 Introduction
Figure 1: Two images containing examples of “things that recur" (i.e. RPs). Due to a lack of explicit straight lines, automatic vanishing point detection on these images poses challenges to explicit line-based methods. The VPD results shown here are from R-VPD (our method), NeurVPS [46], GPVPD [25] and J-Linkage [42] respectively, where the GPVPD and J-Linkage’s VP predictions are too far from the ground truth VP ($\times$) to be displayed here. The top row represents ground truth and the bottom row represents different VP predictions.
@@ 3.1 Implicit Recurring Correspondences Discovery
Figure 3: A Unit Recurring Pattern (URP) is defined to be an RP that has at least two RP instances (RPIs), each of which has at least two distinct visual words (e.g. primitive features such as SIFT) that correspond across the RPIs (figure adapted from [44] ). Here we add the two blue dotted lines passing through pairs of corresponding features, $(f_{11},f_{12})$ and $(f_{21},f_{22})$ , whose intersections identify the VP.
@@ 3.4 Implicit Line Fitting and Weighted RANSAC
Table 1: RPVP-Synthetic dataset performance measured by $AA^{\theta^{\circ}}$ representing the AUC of angle accuracy at the threshold angle $\theta^{\circ}$. The p-values are computed against R-VPD (Ours) to indicate whether the difference in performance is statistically significant. Corresponding AA curves are shown in Figure 6.
@@
Table 2: RPVP-Real dataset performance measured by $AA^{\theta^{\circ}}$ and p-value. Corresponding AA curves are shown in Figure 7.
@@ 4.1.1 RPVP-Synthetic: Recurring Pattern-based Vanishing Point Dataset with Synthetic Images
We created a custom dataset using Blender, an open-source 3D computer graphics software tool [1]. We generate 3,200 images from 16 different synthetic 3D scenes (shown in Figure 4), each contributing 200 images. In this dataset, objects are placed along a straight line and oriented toward a vanishing point, with each scene featuring varying objects, backgrounds, scales, and camera orientations. Ground truth annotations including the $(x,y)$ coordinates of the vanishing point and the camera parameters are provided for each image. More information on scene creation and frame extraction is provided in the Appendix. It is worth noting that a majority of images from this dataset do not contain explicit straight lines oriented towards the vanishing point.
@@ 4.1.3 TMM17-Test: Natural Scenes Dataset
The TMM17 dataset [48, 46] consists of natural scenes with one dominant vanishing point, most containing edges or straight lines oriented towards the VP. We use the test image set of 275 images from [46] (provided through their training (2000 images) and validation (275 images) text files), TMM17-Test, as a validation image set for all VPD methods. Even though our proposed method R-VPD is not designed to deal with images without RPs, we carry out validation on TMM17-Test to get a sense of the robustness of R-VPD on real world scenes with or without recurring patterns.
@@ 4.1.4 RPVP-Real-Exclusive
The RPVP-Real dataset has an overlap of about 70% with the TMM17 dataset (Sec 4.1.3) used to train NeurVPS-tmm17 [46]. To avoid potential data leaking when comparing with NeurVPS, we form an additional dataset which is the subset (30%) of RPVP-Real that has no overlap with TMM17, which we call RPVP-Real Exclusive dataset that consists of 416 images (Figure 5).
Figure 5: The relationship between TMM17, RPVP-Real, and RPVP-Real Exclusive. TMM17 contains generic natural images, while RPVP-Real contains images with RPs only.
@@ 5.2 RPVP-Real-Exclusive: RP Image Exclusive
Table 3: Performance of NeurVPS and R-VPD on RPVP-Real Exclusive dataset (Figure 8).
@@
Table 4: Performance of all VPD methods on TMM17-Test set. Corresponding AA curves are shown in Figure 9.
@@ 5.3 TMM17-Test: Going beyond RP images
To evaluate on images that do not necessarily have recurring patterns in them, we perform an evaluation on the same test set used in [46], TMM17-Test, to compare all VPD methods here. Figure 9 and Table 4 provide AUC for all the methods on TMM17-Test. R-VPD performs second to the best on TMM17-Test without any training, and with a p-value of $p=0.003$ ($>0.001$) the best performer NeurVPS-tmm17 does not significantly outperform our approach.
@@ B.2.1 RPVP-Real Dataset Files:
Figure 15: Robustness of our vanishing point detection system is evaluated using a stress test. We methodically vary the feature densities by resizing images to simulate different levels of feature availability and manipulate spatial distributions by introducing Gaussian noise at various standard deviations ([0.5, 1, 2, 3, 5]) to the location of the feature points. F1-scores are computed at each noise level.
@@ Appendix D Stress Test
To emphasize the robustness of our vanishing point detection system, we conducted a controlled stress test [26]. This test methodically varies feature densities by resizing images to simulate different levels of feature availability and manipulates spatial distributions by introducing Gaussian noise at various standard deviations ([0.5, 1, 2, 3, 5]) to the location of the feature points. We calculated the F1-score for each noise level at angle thresholds $5^{\circ}$ and $10^{\circ}$. The consistency between precision and recall at each threshold simplified our analysis, allowing us to use F1-scores to evaluate performance directly. Figure 15 represents F1-scores calculated at the aforementioned noise levels at angle thresholds $5^{\circ}$ and $10^{\circ}$. These robust results validate the effectiveness of our approach when faced with significant alterations in feature density and spatial arrangement, confirming its applicability across diverse operational environments.
@@ Appendix E More Qualitative Examples
Figure 18: Failure cases of our method compared to other methods. The images are taken from the TMM17-test dataset. Failures are mostly attributed to the absence of RPs and straight-line segments in the image.

### v2

@@ 1 Introduction
Figure 1: Two images containing examples of “things that recur" (Recurring Patterns). Top row: images with ground truth VP ($\times$) indicated. Bottom row: VP prediction results from four VPD methods R-VPD (our method), NeurVPS [46], GPVPD [25] and J-Linkage [42] respectively, where VP predictions that are too far from the ground truth are not shown in the images ((a) 2 misses, (b) 3 misses). Due to a lack of explicit straight lines, automatic vanishing point detection on these images poses challenges to explicit line-based methods.
@@ 3.1 Implicit Recurring Correspondences Discovery
Figure 3: A Unit Recurring Pattern (URP) is an RP with at least two RP instances (RPIs) and each RPI has at least two distinct visual words (e.g. clustered SIFT descriptors) that correspond across the RPIs (this figure is adapted from [44] ). Here we add the two blue dotted lines passing through pairs of corresponding features, $(f_{11},f_{12})$ and $(f_{21},f_{22})$ , whose intersections identify the VP.
@@ 3.4 Implicit Line Fitting and Weighted RANSAC
Table 1: RPVP-Synthetic dataset (3200 images) performance measured by $AA^{\theta^{\circ}}$ representing the AUC of angle accuracy at the threshold angle $\theta^{\circ}$. The p-values are computed against R-VPD (Ours) to indicate whether the difference in performance is statistically significant. Corresponding AA curves are shown in Figure 6.
@@
Table 2: RPVP-Real dataset (1400 images) performance measured by $AA^{\theta^{\circ}}$ and p-value. Corresponding AA curves are shown in Figure 7.
@@ 4.1.1 RPVP-Synthetic: Recurring Pattern-based Vanishing Point Dataset with Synthetic Images
A custom dataset is created using Blender, an open-source 3D computer graphics software tool [1]. We generate 3,200 images from 16 different synthetic 3D scenes (shown in Figure 4), each contributing 200 images. In this dataset, objects are placed along a straight line and oriented toward a vanishing point, with each scene featuring varying objects, backgrounds, scales, and camera orientations. Ground truth annotations including the $(x,y)$ coordinates of the vanishing point and the camera parameters are provided for each image. More information on scene creation and frame extraction is provided in the Appendix. It is worth noting that a majority of images from this dataset do not contain explicit straight lines oriented towards the vanishing point.
@@ 4.1.3 TMM17-Test: Natural Scenes Dataset
The TMM17 dataset [48, 46] consists of 2275 images of natural scenes with one dominant vanishing point. As illustrated in Figure 5 TMM17 overlaps with over 70% of RPVP-Real images (images with recurring patterns). The one split[46] used 2000 images for training and 275 images as the test image set. To avoid any data leaking, we choose TMM17-Test, as a validation image set for all VPD methods. It is worth noting that TMM17-Test contains both images with and without RPs (Figure 5). Even though R-VPD is designed primarily to deal with images with RPs, we use TMM17-Test to test the robustness of R-VPD on real world scenes beyond those with recurring patterns.
@@ 4.1.4 RPVP-Real-Exclusive
The RPVP-Real dataset has an overlap of about 70.3% with the TMM17 dataset (Sec 4.1.3) used to train NeurVPS-tmm17 [46]. To avoid potential data leaking when comparing with NeurVPS, we form an additional dataset which is the subset (29.7%) of RPVP-Real that has no overlap with TMM17, which we call RPVP-Real Exclusive dataset that consists of 416 images (Figure 5).
Figure 5: The relationship between RPVP-Real, TMM17, RPVP-Real Exclusive and TMM17-Test. TMM17/TMM17-Test contains natural images with/without RPs while RPVP-Real/RPVP-Real Exclusive contains images with RPs only.
@@ 5.2 RPVP-Real-Exclusive: RP Image Exclusive
Table 3: Performance of NeurVPS and R-VPD on RPVP-Real Exclusive dataset (416 images). Corresponding AA curves are shown in Figure 8.
@@
Table 4: Performance of all VPD methods on TMM17-Test set (275 images, Fig. 5). Corresponding AA curves are shown in Fig 9.
@@ 5.3 TMM17-Test: Going beyond RP images
To evaluate on images that do not necessarily have recurring patterns in them, we perform an evaluation on the same test set used in [46], TMM17-Test.Figure 9 and Table 4 provide AUC for all the methods on TMM17-Test. R-VPD performs second to the best on TMM17-Test without any training, and with a p-value of $p=0.003$ ($>0.001$) the best performer NeurVPS-tmm17 does not significantly outperform our approach. It is worth noting that of the 275 images in the TMM17-Test set, a considerable portion has no RP on the images (Figure 5).
@@ B.2.1 RPVP-Real Dataset Files:
Figure 15: Robustness of our vanishing point detection system is evaluated using a stress test. We methodically vary the feature densities by resizing images to simulate different levels of feature availability and manipulate spatial distributions by introducing Gaussian noise at various standard deviations ([0.5, 1, 2, 3, 5]) to the location of the feature points. F1-scores are computed at each noise level. Figure 16 shows some sample images.
@@ Appendix D Stress Test
To emphasize the robustness of our vanishing point detection system, we conducted a controlled stress test [26]. This test methodically varies feature densities by resizing images to simulate different levels of feature availability and manipulates spatial distributions by introducing Gaussian noise at various standard deviations ([0.5, 1, 2, 3, 5]) to the location of the feature points. We calculated the F1-score for each noise level at angle thresholds $5^{\circ}$ and $10^{\circ}$. The consistency between precision and recall at each threshold simplified our analysis, allowing us to use F1-scores to evaluate performance directly. Figure 15 represents F1-scores calculated at the aforementioned noise levels at angle thresholds $5^{\circ}$ and $10^{\circ}$. Figure 16 represents the variations in the input image. The top row represents scaled input images that vary the feature density in the input image and the bottom row represents images where noise is added to the location of the SIFT features to manipulate spatial distribution. It can be seen that the SIFT features in the rightmost image (bottom row) with $std=5$ is more spread out compared to the middle image (bottom row) with $std=0.5$. These robust results validate the effectiveness of our approach when faced with significant alterations in feature density and spatial arrangement, confirming its applicability across diverse operational environments.
Figure 16: Top row: original image and scaled images that introduce density variation of SIFT features. Bottom row: zoomed-in image patches representing SIFT features with noise added to their location that perturbs the spatial distribution of the features.
@@ Appendix E More Qualitative Examples
Figure 19: Failure cases of our method. The images are all from the TMM17-test dataset (Figure 5). Failures are mostly attributed to the absence of RPs as well as straight-line segments in the image.